# Hallucinations and Key Information Extraction in Medical Texts: A Comprehensive Assessment of Open-Source Large Language Models

### v1

@@ Abstract
Clinical summarization is crucial in healthcare as it distills complex medical data into digestible information, enhancing patient understanding and care management. Large language models (LLMs) have shown significant potential in automating and improving the accuracy of such summarizations due to their advanced natural language understanding capabilities. These models are particularly applicable in the context of summarizing medical/clinical texts, where precise and concise information transfer is essential. In this paper, we investigate the effectiveness of open-source LLMs in extracting key events from discharge reports, such as reasons for hospital admission, significant in-hospital events, and critical follow-up actions. In addition, we also assess the prevalence of various types of hallucinations in the summaries produced by these models. Detecting hallucinations is vital as it directly influences the reliability of the information, potentially affecting patient care and treatment outcomes. We conduct comprehensive numerical simulations to rigorously evaluate the performance of these models, further probing the accuracy and fidelity of the extracted content in clinical summarization.
@@
Medical Text Summarization, Large Language Models, Hallucinations, Key Information Extraction, Factual Accuracy.
@@ I Introduction
Large Language Models (LLMs) have emerged as powerful tools in artificial intelligence and machine learning, demonstrating remarkable capabilities in natural language understanding, generation, and contextual reasoning [6]. These models, trained on vast and diverse datasets, can comprehend and generate human-like text, making them particularly well-suited for tasks such as summarization, translation, and question-answering [7]. In the medical domain, LLMs offer significant potential for summarizing clinical texts [8] by capturing key information while preserving critical details necessary for decision-making. Unlike traditional rule-based or statistical summarization methods, LLMs can adapt to complex linguistic structures, recognize medical terminologies, and generate concise yet informative summaries tailored to specific clinical needs. By leveraging contextual embeddings and domain-specific fine-tuning, LLMs can improve the accuracy and relevance of medical text summarization, ultimately supporting healthcare professionals in reducing information overload and enhancing patient care [6].
@@
Hallucinations: LLMs are known to generate plausible yet incorrect or non-existent information, a phenomenon referred to as hallucination [9, 10]. In the medical domain, hallucinated details in summary, such as incorrect dosages, fabricated conditions, or misrepresented patient histories, thus can have serious consequences for clinical decision-making. This challenge underscores the need for robust validation mechanisms, domain-specific fine-tuning, and external fact-checking to ensure the reliability of medical text summaries [11].
In this paper, we aim to highlight and analyze the challenges of key event extraction and hallucinations in LLM-driven medical text summarization. To achieve this, we conduct a comprehensive evaluation using discharge summaries from the MIMIC-IV dataset [12], a widely used repository of de-identified clinical records. We explore various open-source large language models (LLMs) to assess their effectiveness in extracting essential medical events while minimizing hallucinations. The open-source LLMs used in this study include LLaMA [13], Mistral [14], Gemma [15], Phi [16], Falcon [17], LLaVA [18], DeepSeek [19], and Qwen [20]. Our analysis focuses on understanding the strengths and limitations of these models in handling domain-specific language and preserving critical clinical information. Addressing these issues is essential to making LLM-driven summarization a practical and trustworthy tool for healthcare applications.
@@ II-A1 Key Event/Information Extraction
With the advent of LLMs, researchers have explored their potential for key event extraction in clinical text summarization. Models such as GPT-4 [25] and Med-PaLM [26] have shown promise in capturing contextual dependencies, but they may still fail to prioritize clinically relevant insights accurately. Furthermore, without explicit fine-tuning on labeled clinical datasets, LLMs risk omitting critical patient information, making their direct application to medical summarization challenging.
@@ II-A2 Detection of Hallucination
Existing research has explored various techniques for hallucination detection and mitigation in NLP applications. Some approaches rely on fact-checking methods that compare generated summaries against source texts using similarity metrics or retrieval-based validation. Others employ uncertainty quantification techniques, such as confidence scoring and probabilistic modeling, to assess the reliability of LLM-generated outputs. In the biomedical domain, external knowledge bases (e.g., SNOMED CT [27]) have been integrated into NLP pipelines to verify the factual accuracy of generated content. However, ensuring faithfulness in LLM-generated medical summaries remains an open challenge, necessitating further research into robust evaluation metrics and hallucination filtering mechanisms.
@@ III Analyzing LLM-Generated Clinical Summaries
In this section, we examine the key aspects of essential information extraction and corresponding hallucinations in clinical summarization by LLMs. To illustrate these, we provide a detailed analysis using a representative hospital discharge report as an example.
@@ III-A Key Event Extraction in Clinical Summarization
Text summarization is a key task in natural language processing, condensing extensive texts into concise summaries while preserving essential information. It is widely used across domains requiring swift data comprehension, such as finance, where it streamlines market reports and investment analyses by highlighting critical trends and risks, and law, where it simplifies legal documents to emphasize case facts, arguments, and verdicts. In medical applications, summarization extracts vital clinical details from documents like patient histories and treatment plans, aiding both healthcare professionals and patients. For example, summarizing a CT scan report can highlight key findings, facilitating swift clinical decisions. Leveraging open-source LLMs for medical text summarization enhances patient care by improving communication and clarity of critical information.
@@ III-B Hallucination in LLM-Generated Summaries
Firstly, the Qwen2.5 summary inaccurately assigns the patient an age of $45$, a detail absent from the original report, exemplifying an unsupported fact hallucination. Additionally, it omits essential follow-up instructions for Liver Function Tests (LFTs), despite the report explicitly stating their necessity twice. With 122 characters still available, this omission reflects a faithfulness hallucination, where the model fails to capture critical details vital for ongoing patient care.
@@ IV-A Open-source LLMs and Dataset
In this study, we utilize open-source large language models (LLMs) to summarize hospital discharge reports. The selected LLMs include LLaMA, DeepSeek, and Phi etc. Table I provides an overview of these models along with their respective parameter counts. Here, we employ a subset of hospital discharge reports sourced from the publicly available MIMIC-IV dataset [29]. Each discharge report is sequentially processed using each LLM’s standardized prompt.
@@ IV-B Key Information Extraction
LLM # Parameters Avg. Length (SD) “Comprehensively” Covered “Fairly” Covered
Reasons Events Follow-up Reasons Events Follow-up
LLaMA3.1 $8B$ $674(95)$ $70.33\%$ $58\%$ $29.33\%$ $71.67\%$ $59.67\%$ $30.33\%$
Qwen2.5 $7B$ $778(160)$ $83.33\%$ $70\%$ $50.67\%$ $85\%$ $70.67\%$ $52.33\%$
DeepSeek-v2 $16B$ $1010(332)$ $80.33\%$ $65\%$ $41.33\%$ $81\%$ $67.33\%$ $42.67\%$
Phi3 $3.8B$ $1695(580)$ $81.33\%$ $63\%$ $55\%$ $83\%$ $65.67\%$ $58.33\%$
Gemma2 $9B$ $621(124)$ $69.67\%$ $55.67\%$ $30\%$ $70.67\%$ $58.33\%$ $31.33\%$
MistralLite $7B$ $1400(795)$ $77.67\%$ $69\%$ $49\%$ $78.67\%$ $70\%$ $50\%$
LLaVA $7B$ $993(416)$ $70.67\%$ $61.67\%$ $53.33\%$ $73.33\%$ $64.33\%$ $57.33\%$
We assess open-source LLMs on extracting (i) admission reasons, (ii) key events, and (iii) follow-ups from discharge reports, using GPT-4 as an evaluator [28]. Table I shows the percentage of key details captured, highlighting each model’s strengths and limitations in medical text understanding.
@@
From the results, Qwen2.5 and DeepSeek-v2 emerge as the top-performing models in extracting admission reasons, with Qwen2.5 achieving $83.33\%$ comprehensively and $85\%$ fairly. Similarly, Phi3 and DeepSeek-v2 show strong performance in capturing hospitalization events, while LLaMA3.1 and LLaVA exhibit relatively lower effectiveness. Follow-up recommendations prove to be the most challenging category for all models, with comprehensive coverage ranging from $29.33\%$ (LLaMA3.1) to $55\%$ (Phi3). However, fair coverage is comparatively better, with Phi3 and LLaVA achieving above $57\%$. These indicate that while LLMs perform reasonably well in extracting reasons for admission and key hospitalization events, their effectiveness in summarizing necessary follow-ups remains limited and requires further enhancement.
@@
The observed limitations in extracting key events cannot be solely attributed to the upper bound of $1000$ characters set for the summaries. As shown in Table I, many of the generated summaries fall well below this limit, yet they still fail to capture essential details such as reasons for admission, key hospitalization events, and necessary follow-ups. This suggests that the issue lies more in how the models prioritize information rather than in the imposed length constraint. Interestingly, some models, such as Phi3 and MistralLite, exceeded the given character limit, indicating that the restriction was not strictly enforced across all models. Their performance could improve with fine-tuning to better use available space and ensure comprehensive event extraction.
@@ IV-C Hallucinations
The results reveal notable variations in hallucination tendencies among models. Phi3 exhibits the highest number of hallucinations in both categories, with 150 unsupported fact hallucinations and 111 incorrect/contradicted fact hallucinations, indicating significant reliability concerns. DeepSeek-v2, while generating fewer unsupported facts (46), shows a high number of contradicted facts (102), suggesting issues in accurately interpreting medical information. Other models, such as LLaMA3.1, Qwen2.5, and Gemma2, show more balanced yet still concerning levels of hallucinations, while MistralLite and LLaVA tend to generate a higher number of unsupported facts. These findings highlight the need for better fine-tuning and fact-verification to improve the factual reliability of LLM-generated medical summaries in clinical use.
@@ V Conclusion
Our study of open-source large language models (LLMs) for summarizing hospital discharge reports reveals their potential to accurately extract crucial clinical information but also highlights their susceptibility to generating hallucinations, including unsupported and incorrect facts. These inaccuracies pose risks to patient care and safety. Our comprehensive numerical simulations indicate variability in model performance, underscoring the need for model refinement to reduce hallucinations and enhance understanding of medical contexts. Future work can focus on fine-tuning models to enhance summary quality and mitigate hallucinations. Additionally, incorporating robust validation methods within clinical workflows can improve the reliability of LLM-generated summaries, providing healthcare professionals with efficient and accurate tools for patient care.

### v3

@@ Abstract
Clinical summarization is crucial in healthcare as it distills complex medical data into digestible information, enhancing patient understanding and care management. Large language models (LLMs) have shown significant potential in automating and improving the accuracy of such summarizations due to their advanced natural language understanding capabilities. These models are particularly applicable in the context of summarizing medical/clinical texts, where precise and concise information transfer is essential. In this paper, we investigate the effectiveness of open-source LLMs in extracting key events from discharge reports, including admission reasons, major in-hospital events, and critical follow-up actions. In addition, we also assess the prevalence of various types of hallucinations in the summaries produced by these models. Detecting hallucinations is vital as it directly influences the reliability of the information, potentially affecting patient care and treatment outcomes. We conduct comprehensive simulations to rigorously evaluate the performance of these models, further probing the accuracy and fidelity of the extracted content in clinical summarization. Our results reveal that while the LLMs (e.g., Qwen2.5 and DeepSeek-v2) perform quite well in capturing admission reasons and hospitalization events, they are generally less consistent when it comes to identifying follow-up recommendations, highlighting broader challenges in leveraging LLMs for comprehensive summarization.
@@
Medical Text Summarization, Large Language Models, Hallucinations, Key Information Extraction.
@@ I Introduction
Large Language Models (LLMs) have emerged as powerful tools in artificial intelligence and machine learning, demonstrating remarkable capabilities in natural language understanding, medical imaging [6], generation, and contextual reasoning [7]. These models, trained on vast and diverse datasets, can comprehend and generate human-like text, making them particularly well-suited for tasks such as summarization, translation, and question-answering [8]. In the medical domain, LLMs offer significant potential for summarizing clinical texts [9] by capturing key information while preserving critical details necessary for decision-making. Unlike traditional rule-based or statistical summarization methods, LLMs can adapt to complex linguistic structures, recognize medical terminologies, and generate concise yet informative summaries tailored to specific clinical needs. By leveraging contextual embeddings and domain-specific fine-tuning, LLMs can improve the accuracy and relevance of medical text summarization, ultimately helping reduce information overload and improve patient care [7].
@@
Hallucinations: LLMs are known to generate plausible yet incorrect or non-existent information, a phenomenon referred to as hallucination [10, 11, 12]. In the medical domain, hallucinated details in summary, such as incorrect dosages, fabricated conditions, or misrepresented patient histories, thus can have serious consequences for clinical decision-making. Fig. 1 illustrates an example where the LLM-generated summary introduces unsupported or fabricated clinical information. It includes hallucinated clinical events (e.g., “heart attack is unlikely”) and incorrect medication details (e.g., “Metformin to 1000 mg”), which were not present in the original discharge report. This issue underscores the need for domain-specific fine-tuning and external fact-checking to ensure the reliability of medical text summaries [13, 14].
In this paper, we aim to highlight and analyze the challenges of key event extraction and hallucinations in LLM-driven medical text summarization. To achieve this, we conduct a comprehensive evaluation using discharge summaries from the MIMIC-IV dataset [15], a widely used repository of de-identified clinical records. We explore various open-source large language models (LLMs) to assess their effectiveness in extracting essential medical events while minimizing hallucinations. The open-source LLMs used in this study include LLaMA [16], Mistral [17], Gemma [18], Phi [19], Falcon [20], LLaVA [21], DeepSeek [22], and Qwen [23]. Our analysis examines the strengths and limitations of these models in handling domain-specific language and preserving critical clinical information, which is essential for making LLM-driven summarization a practical and trustworthy tool for healthcare applications.
@@ II-A1 Key Event/Information Extraction
With the advent of LLMs, researchers have explored their potential for key event extraction in clinical text summarization. Models such as GPT-4 [28] and Med-PaLM [29] have shown promise in capturing contextual dependencies, but they may still fail to prioritize clinically relevant insights accurately. Furthermore, without explicit fine-tuning on labeled clinical datasets, LLMs risk omitting critical patient information, making their direct use to medical summarization challenging.
@@ II-A2 Detection of Hallucination
Existing research has explored various techniques for hallucination detection and mitigation in NLP applications. Some approaches rely on fact-checking methods that compare generated summaries against source texts using similarity metrics or retrieval-based validation. Others employ uncertainty quantification techniques, such as confidence scoring and probabilistic modeling, to assess the reliability of LLM-generated outputs. In the biomedical domain, external knowledge bases (e.g., SNOMED CT [30]) have been integrated into NLP pipelines to verify the factual accuracy of generated content. However, ensuring faithfulness in LLM-generated medical summaries remains an open challenge, necessitating further research into robust evaluation metrics and hallucination filters.
@@ III Analyzing LLM-Generated Clinical Summaries
In this section, we examine the key aspects of essential information extraction and corresponding hallucinations in clinical summarization by LLMs. We illustrate these through a detailed analysis of a representative hospital discharge report, demonstrating both accurate extractions and common errors made by LLMs in real-world clinical settings.
@@ III-A Key Event Extraction in Clinical Summarization
Text summarization is a key task in natural language processing, condensing extensive texts into concise summaries while preserving essential information. It is widely used in domains requiring swift data comprehension, e.g., finance and law, to simplify legal documents while highlighting key trends, risks, facts, and arguments. In medical applications, summarization extracts vital clinical details from documents like patient histories and treatment plans, aiding both healthcare professionals and patients. For example, summarizing a CT scan report can highlight key findings, facilitating swift clinical decisions. Leveraging open-source LLMs for medical text summarization enhances patient care by improving communication and clarity of critical information.
@@ III-B Hallucination in LLM-Generated Summaries
Firstly, the Qwen2.5 summary incorrectly states the patient is $45$ years old: an unsupported fact not found in the report. It also omits essential follow-up instructions for Liver Function Tests (LFTs), despite the report explicitly stating their necessity twice. This is particularly concerning given that timely LFT monitoring can be crucial for managing the patient’s complications noted during hospitalization. With 122 characters still available, this omission reflects a faithfulness hallucination, where the model fails to include essential care details.
@@
Our evaluation method, focused on extracting key information and spotting mistakes, provides useful insights for creating better discharge summaries. These results can help clinical teams improve language model tools that support usability, readability, and patient comprehension in real-world healthcare settings by by ensuring accurate and clear communication.
LLM # Parameters Avg. Summary Length (SD) “Comprehensively” Covered “Fairly” Covered
Reasons Events Follow-up Reasons Events Follow-up
LLaMA3.1 $8B$ 674(95)674\;(95) $70.33\%$ $58\%$ $29.33\%$ $71.67\%$ $59.67\%$ $30.33\%$
Qwen2.5 $7B$ 778(160)778\;(160) $\mathbf{83.33\%}$ $\mathbf{70\%}$ $50.67\%$ $\mathbf{85\%}$ $\mathbf{70.67\%}$ $52.33\%$
DeepSeek-v2 $16B$ 1010(332)1010\;(332) $80.33\%$ $65\%$ $41.33\%$ $81\%$ $67.33\%$ $42.67\%$
Phi3 $3.8B$ 1695(580)1695\;(580) $81.33\%$ $63\%$ $\mathbf{55\%}$ $\mathbf{83\%}$ $65.67\%$ $\mathbf{58.33\%}$
Gemma2 $9B$ 621(124)621\;(124) $69.67\%$ $55.67\%$ $30\%$ $70.67\%$ $58.33\%$ $31.33\%$
MistralLite $7B$ 1400(795)1400\;(795) $77.67\%$ $69\%$ $49\%$ $78.67\%$ $70\%$ $50\%$
LLaVA $7B$ 993(416)993\;(416) $70.67\%$ $61.67\%$ $53.33\%$ $73.33\%$ $64.33\%$ $57.33\%$
@@ IV-A Open-source LLMs and Dataset
In this study, we utilize open-source LLMs to summarize hospital discharge reports. We selected a diverse set of LLMs for evaluation, as discussed in Sec. I. Table I provides an overview of these models along with their respective parameter counts. Here, we employ a subset of 100 hospital discharge reports sourced from the publicly available MIMIC-IV dataset, consistent with those used in [32]. Each discharge report is sequentially processed using each LLM’s standardized prompt.
@@ IV-B Key Information Extraction
We assess open-source LLMs on extracting (i) hospital admission reasons, (ii) key events occurred during the hospital stay, and (iii) follow-ups, from the hospital discharge reports, using GPT-4 as an evaluator [31]. The key information extraction step uses the following prompt:
From the provided hospital discharge report, identify (i) three primary reasons for the patient’s admission, (ii) three most important events that occurred during the hospital stay, and (iii) three crucial follow-up actions required after the discharge from the hospital.
Next, we evaluate whether these extracted elements are reflected in the LLM-generated summary. The following prompt is used:
Given the extracted hospital information: admission reasons, hospital events and follow-up actions; and the provided model-generated summary: “the corresponding summary”, identify how many of the reasons, important events, and follow-up actions are fairly covered in the summary.
Note that we acknowledge the potential risk of evaluator bias when using GPT-4 [35], however, we selected GPT-4 due to its high agreement with domain experts in prior studies on medical summarization. We also manually reviewed a subset of samples to calibrate model judgments and found strong alignment with human expectations. Figure 4 shows an original hospital discharge report and the corresponding LLaMA-generated summary. In this example, GPT-4 correctly identifies that, while LLaMA captures the admission reasons well, it falls short in covering the follow-up actions. Next, Table I shows the percentage of key details captured, highlighting each model’s strengths and limitations in medical text understanding. The corresponding outputs from open-source LLMs and the judge model (GPT) are available in [36].
[…]
Summary by LLaMA: You were admitted to the hospital due to a clotted AV (arteriovenous) fistula in your left arm, which was preventing dialysis treatments for kidney failure. Your doctors performed a successful thrombectomy (a procedure to clear the blockage) and you underwent dialysis afterwards without any complications. We also investigated some low blood cell counts, known as pancytopenia, but couldn’t determine the cause during your hospital stay. We will follow up on this with your primary care doctor when you are discharged. Your hepatitis C infection is a chronic condition that we recommended further treatment for, especially since it may have contributed to a previous episode of encephalopathy (brain damage).
Figure 4: Comparison of the original hospital course note and LLaMA-generated summary: while LLaMA summary captures the primary admission reasons (including thrombosed AV fistula and kidney failure) and the hospital events (including thrombectomy and dialysis), it omits key follow-up plans: outpatient discussion for HCV treatment and routine AV graft maintenance every three months.
@@
From the results, Qwen2.5 and DeepSeek-v2 emerge as the top-performing models in extracting admission reasons, with Qwen2.5 achieving $83.33\%$ comprehensively and $85\%$ fairly. Similarly, Phi3 and DeepSeek-v2 show strong performance in capturing hospitalization events, while LLaMA3.1 and LLaVA exhibit relatively lower effectiveness. Follow-up recommendations prove to be the most challenging category for all models, with comprehensive coverage ranging from $29.33\%$ (LLaMA3.1) to $55\%$ (Phi3). A possible explanation is that admission reasons are typically stated early in discharge reports, while follow-up instructions are scattered across different sections, making them harder to capture. However, fair coverage is comparatively better, with Phi3 and LLaVA achieving above $57\%$. These indicate that while LLMs perform reasonably well in extracting reasons for admission and key hospitalization events, their effectiveness in summarizing necessary follow-ups remains limited and requires further enhancement.
@@
The observed limitations in extracting key events cannot be solely attributed to the 1000-character summary limit. As shown in Table I, many summaries fall well below this threshold yet still miss critical details like admission reasons, major events, and follow-ups. This suggests that the issue lies more in how the models prioritize information rather than in the imposed length constraint. Interestingly, some models, such as Phi3 and MistralLite, exceeded the given character limit, indicating that the restriction was not strictly enforced across all models. Their performance could improve with fine-tuning to better use available space and ensure comprehensive event extraction.
@@ IV-C Hallucinations
The results reveal notable variations in hallucination tendencies among models. Phi3 exhibits the highest number of hallucinations in both categories, with 150 unsupported fact hallucinations and 111 incorrect/contradicted fact hallucinations, indicating significant reliability concerns. DeepSeek-v2, while generating fewer unsupported facts (46), shows a high number of contradicted facts (102), suggesting issues in accurately interpreting medical information. Other models, such as LLaMA3.1, Qwen2.5, and Gemma2, show more balanced yet still concerning levels of hallucinations, while MistralLite and LLaVA tend to generate a higher number of unsupported facts. These findings underscore the need for better fine-tuning and fact-verification, positioning this work as a crucial diagnostic step toward safer clinical NLP deployment.
IV-D Evaluator Robustness: Extending to Gemini
While GPT-4 has been used as the primary evaluator throughout this paper, we now assess the reliability of our evaluation framework by incorporating Gemini as an alternative evaluator. Table III presents the results based on Gemini’s assessments. We find that open-source LLMs identify admission reasons quite well with very high accuracy, but often struggle with follow-up actions, with coverage dropping as low as $32\%$. Despite minor variations in scores, the trends and model rankings remain largely consistent with those obtained using GPT-4, suggesting that our findings are robust across strong evaluators. These similarities suggest that our findings are not heavily dependent on the choice of evaluator, reinforcing the validity of our conclusions regarding LLM performance in clinical summarization tasks.
TABLE III: Comparison of LLMs in terms of the total number of extracted key events from the respective summaries using Gemini as the evaluator. Note that we omit Gemma2 from this comparison, as both Gemini and Gemma2 are developed by the same organization, which may introduce additional bias in the evaluation.
LLM Comprehensively Covered
Reasons Events Follow up
LLaMA3.1 $82.00\%$ $69.00\%$ $32.33\%$
Qwen2.5 $86.67\%$ $77.00\%$ $52.33\%$
DeepSeekv2 $91.00\%$ $70.33\%$ $44.67\%$
Phi3 $81.67\%$ $69.33\%$ $57.00\%$
MistralLite $75.00\%$ $64.67\%$ $44.67\%$
LLaVA $76.67\%$ $63.00\%$ $48.67\%$
@@ V Conclusion
Our study of open-source large language models (LLMs) for summarizing hospital discharge reports reveals their potential to accurately extract crucial clinical information but also highlights their susceptibility to generating hallucinations, including unsupported and incorrect facts. These inaccuracies pose risks to patient care and safety. Our simulations show performance variability: while models (e.g. Qwen2.5 and DeepSeek-v2) capture admission reasons and events well, they often miss follow-up recommendations, highlighting the need for refinement and better medical understanding.
Future work can focus on developing automatic hallucination detection techniques [37] to identify and flag inaccurate or unsupported content in clinical text summaries. Fine-tuning models to enhance summary quality and reduce hallucinations can be another key direction. Additionally, integrating robust validation methods into clinical workflows can improve the reliability of LLM-generated summaries, offering healthcare professionals more efficient and accurate tools for patient care. Evaluating these summaries using readability metrics (e.g., Flesch-Kincaid) can further help assess their clarity and usefulness for patients. The analysis could be extended beyond discharge summaries to include other clinical documents such as progress and radiology reports. Furthermore, one can explore domain-specific hallucination clustering and temporal reasoning challenges in extracting follow-up instructions, aiming to guide safer real-world deployment of clinical NLP systems.